# KeyDiff: Key Similarity-Based KV Cache Eviction for Long-Context LLM Inference in Resource-Constrained Environments

### v2

@@ 3.3 Derivation of KeyDiff from an Optimization Perspective
Figure 4: (a, b, and c) PCA Visualizations in two dimensions of a key cache managed with Sink, TOVA, and KeyDiff. Retained tokens are blue, while evicted tokens are orange. The retained tokens using Sink and TOVA cluster together, while KeyDiff retains more diverse keys. Keys are taken from layer $5$ and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset. (d) PCA visualization of the retained keys for each KV cache eviction method. We present distributions of additional heads and layers in Section A.3.
4 Experiment
In this section, we empirically demonstrate the effectiveness of KeyDiff. We begin with a description of competing, state-of-the-art eviction methods, followed by a detailed description of the evaluation setup. We then test each eviction policy on the LongBench task suite (Section 4.2) and report results in Table 1 followed by an ablation study (Section 4.3) of important properties of KeyDiff. Our findings can be summarized as follows:
  • KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Kamradt, 2023) (Section 4.1).
  • KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section 4.2).
  • KeyDiff, utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector, generally outperforms other variants of KeyDiff that use different similarity metrics or anchor vectors. (Section 4.3).
@@ 4 Experiment
We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama $3.1$-$8$B-Instruct (Dubey et al., 2024), Llama $3.2$-$3$B-Instruct, and Qwen $2.5$-$3$B/7B-Instruct. We evaluate these models using H2O (Zhang et al., 2024b), TOVA (Oren et al., 2024), SnapKV (Li et al., 2024), and StreamingLLM (Xiao et al., 2024), (or “sink attention”) cache eviction policies, along with the eviction-free model performance as a baseline. We simulate a resource constrained environment by processing prompts and generating responses using Equation 6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments. We denote the cache budgets of 2048, 4096, 6144 and 8192 as 2K, 4K, 6K and 8K, respectively.
4.1 Needle In a Haystack
To compare the impact of various cache eviction policies on fact retrieval, we conduct the “Needle In a Haystack” test (Liu et al., 2024a; Kamradt, 2023). This test embeds specific information (“needle”) at different points within a body of unrelated text (“haystack”); finding and retaining the needle is challenging for eviction policies, which can’t know during eviction what information must be retained. The results are shown in Figure 6, where we show the recall accuracy of Llama3.23.23.23.2-$3$B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of $6$K. KeyDiff performs similarly to TOVA and sink attention for shorter documents and outperforms both methods as the document length increases.
(a) Sink
(b) TOVA
@@ 4.2 LongBench
LongBench (Bai et al., 2024) is a bilingual, multi-task benchmark suite for LLMs, providing a comprehensive stress test for long prompt inputs. LongBench is useful to evaluate cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K. For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section B.3 for more detail.)
Table 1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$. We present the full evaluation results in Table 11, and those of Qwen 2.5-3B/7B-Instruct in Table 9.
As shown in Table 1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets. KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized (Bai et al., 2024), while achieving near full-context model performance even with the smallest budget. We observed similar trends in the full LongBench task suite as shown in Table 11.
Results for Full Prompt Processing
We also evaluate the methods in more standard inference environments without block prompt processing (i.e., setting the block size of the prompt to $B=\infty$) and summarize the results in Table 7. KeyDiff exhibits similar or better performance compared to competing methods. Notably, the attention-based methods (e.g., H2O, TOVA, and SnapKV) show significant performance improvements over the $B=128$ case. This result supports our hypothesis: an eviction scheme robust to changes in the scope of comparison among tokens is essential in memory constrained environments where token-wise attention weight can’t be fully materialized.
@@ 4.3 Ablation Study
We evaluate the design choices of KeyDiff, including the similarity metrics and the choice of the anchor vector, and validate the efficacy of KeyDiff.
Table 2: Ablation study results about anchor vector. Average of Full Longbench results for Llama 3.2-3B-Instruct. KeyDiff results match Table 11. (Higher is better)
2K 4K 6K 8K
KeyDiff 35.09 40.28 41.88 42.40
$\operatorname{Pairwise}$ 35.24 40.61 41.87 42.45
$\operatorname{Median}$ 35.43 40.67 41.89 42.26
Selecting the Anchor Vector
Table 3: Ablation study results about distance metric. Average of Full Longbench results of Llama 3.2-3B-Instruct. KeyDiff results match Table 11. (Higher is better)
2K 4K 6K 8K
KeyDiff 35.09 40.28 41.88 42.40
$\operatorname{DotProd}$ 30.14 38.01 41.09 42.23
$\operatorname{Euclidean}$ 13.68 21.06 25.91 29.53
Selecting the Similarity Metric
@@ 5 Related Work
LLMs often exhibit high attention sparsity, where a small subset of keys receives a significant proportion of attention scores. This characteristic allows sparse approximation techniques to reduce the computational cost of attention. Similar to PagedAttention (Kwon et al., 2023), Tang et al. (2024) estimates the importance of a page (a contiguous set of keys) to a given query, whereas Rehg (2024) further refined the budgets in a per-head manner. On the contrary, sample-based methods (Zhu et al., 2024; Ribar et al., ) attempt to approximate token importance by inspecting the attention scores from the last few queries or certain query channel dimensions. Despite their effectiveness in reducing computational costs, these methods do not address the memory overhead of the KV cache, which typically retains all KVs.
@@
Different approaches to compress the KV cache include architecture modification such as GQA (Ainslie et al., 2023), which shares a KV cache across a small number of heads. Other techniques to compress the KV cache include quantization such as in (Hooper et al., 2024; Liu et al., 2024b; Zhang et al., 2024a) in which the authors use various techniques to take advantage of existing patterns to efficiently quantize and compress the KV cache. More related to our work (Yang et al., 2024b) uses a scoring mechanism to determine the precision of the quantization for different tokens.
@@ 6 Conclusion
Inspired by our observation that distinctive keys tend to have high attention scores, we propose KeyDiff, a training-free KV cache eviction method based on key similarity that enables large language models to operate in memory and compute constrained environments. We justify KeyDiff by showing that it computes the optimal solution for an optimization problem minimizing the pairwise cosine similarity among keys in the KV cache. KeyDiff significantly outperforms state-of-the-art KV cache eviction methods under similar memory constraints, with only a 1.5% and 0.04% accuracy drop from the non-evicting baseline while achieving 33% and 23% KV cache memory reduction on LongBench.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Attention-based eviction methods
  • TOVA (Oren et al., 2024): $\phi^{\mathrm{TOVA}}(A)=A_{-1,:}$,
  • H2O (Zhang et al., 2024b): $\phi^{\mathrm{H2O}}(A)=A_{\mathrm{prev}}+A^{\top}1$,
  • SnapKV (Xiao et al., 2024): $\phi^{\mathrm{SnapKV}}(A)=(A^{\top}1)\ast K$, where $K$ is a vector of $\frac{1}{k}$ and $k$ is the kernel size of average smoothing.
A.2 Runtime and Memory Complexity
@@
A.2.1 FLOP count of KeyDiff
Regarding the FLOP count of KeyDiff: the bulk of the computation in KeyDiff (neglecting the ‘topk‘ operator) is the following two expressions:
@@ A.2.1 FLOP count of KeyDiff
To compute the cosine similarity score, we have from above that $\mu(\hat{K})\cdot k_{i}$ requires $nd$ multiplications and $n(d-1)$ additions. Also from above, we have that computing $\|\mu(\hat{K})\|$ requires $d$ multiplications, $d-1$ additions and one square root. We reuse the computation of $\|k_{i}\|$ from the previous step and compute $\|k_{i}\|\|\mu(\hat{K})\|$ in $n$ multiplications and $\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)$ with more $n$ additions (assuming boolean comparison is equals addition in cost). We can then divide through to compute $\frac{\mu(\hat{K})\cdot k_{i}}{\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)}$ with $n$ divisions. Therefore, computing the cosine similarity between the anchor and each key requires $nd+d+n$ multiplications, $nd+d-1$ additions, $n+1$ divisions and one square root.
@@
If we declare additions and multiplications are roughly one FLOP (i.e. can be computed in one cycle), division is roughly 5 FLOPs and square root is roughly 8 FLOPs, we arrive at a final FLOP count of:
$(3nd+d+n)+(3nd-n-1)+5*(2n+2)+8*(n+1)=(6d+18)n+d+17.$ (A.1)
@@
A.3 Additional PCA Visualizations
@@
Appendix B Additional discussion for LongBench
B.1 Empirical Motivation for KeyDiff Setup

### v3

@@ 3.3 Why KeyDiff Works: A Theoretical Perspective
Figure 5: (a, b, and c) PCA Visualizations in two dimensions of a key cache managed with Sink, TOVA, and KeyDiff. Retained tokens are blue, while evicted tokens are orange. Keys are taken from layer $5$ and head $3$ of Llama3.2-3B-Instruct, and generated using the NarrativeQA dataset. (d) PCA visualization of the retained keys for each KV cache eviction method.
4 Experiments
In this section, we empirically demonstrate the effectiveness of KeyDiff. We begin with a description of competing, state-of-the-art eviction methods, followed by a detailed description of the evaluation setup, then present our experimental results. Our findings can be summarized as follows:
  • Needle-In-a-Haystack. KeyDiff outperforms competing eviction policies on the Needle-In-A-Haystack benchmark (Section 4.1).
  • LongBench. KeyDiff outperforms competing eviction policies with block size $B=128$ on LongBench, achieving an 1.5% accuracy drop with a 6K cache budget ($\sim$33compression rate) and $\leq$ .04% with a 8k cache budget ($\sim$23compression rate) with Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct (Section 4.2).
  • Reasoning. KeyDiff performs competitively on the Math-500 reasoning benchmark with other eviction methods using the DeepSeek-R1-Distill-Qwen-7B and Llama-8B, and shows near eviction-free baseline performance when augmented with a sliding window (Section 4.3) for DeepSeek-R1-Distill-Llama-8B.
  • Ablation Study. We perform an ablation study on the main parameters of KeyDiff and show that utilizing negative cosine similarity as the eviction criteria and the mean of cached keys as the anchor vector performs best. (Section 4.4).
  • Efficiency. We compare the end-to-end inference latency of KeyDiff, [20] and [24] and observe a 30% latency improvement with KeyDiff (Section 4.5).
@@ 4 Experiments
We apply several cache eviction methods to several decoder-only transformer-based language models, including Llama $3.1$-$8$B-Instruct [10], Llama $3.2$-$3$B-Instruct [10], and Qwen $2.5$-$3$B/7B-Instruct [38]. We evaluate these models using H2O [43], TOVA [24], SnapKV [20], and StreamingLLM [35], (or “sink attention") cache eviction policies, along with the eviction-free model as a baseline. We simulate a resource constrained environment by processing prompts and generating responses using Equation 6, with a block size of $B=128$ for prompt processing and $B=1$ for token generation using greedy decoding for all experiments. We denote the cache budgets of 2048, 4096, 6144 and 8192 as 2K, 4K, 6K and 8K, respectively.
(a) TOVA
(b) SnapKV
@@
4.1 Needle In a Haystack
To compare the impact of various cache eviction policies on fact retrieval, we conduct the “Needle In a Haystack" test [21, 16]. This test embeds specific information (“needle") at different points within a body of unrelated text (“haystack"); finding and retaining the needle is challenging for eviction policies, which can’t know what information must be retained during block prompt processing. The results are shown in Figure 6 and Figure 19, where we show the recall accuracy of Llama3.23.23.23.2-$3$B-Instruct across different document lengths (x-axis) and needle depths (y-axis) with a cache size of $6$K. KeyDiff performs similarly to TOVA, SnapKV and sink attention for shorter documents and outperforms all three methods as the document length increases.
@@ 4.2 LongBench
LongBench [4] is a bilingual, multi-task benchmark suite for LLMs, providing a comprehensive stress test for long prompt inputs. LongBench is useful for evaluating cache eviction methods in a resource constrained environments with a fixed memory budget: 51% of prompts are longer than the largest KV cache size of 8K. For cache budgets of 6k and 8k tokens, prompts in LongBench are compressed by 33% and 23% respectively on average, (see Section F.3 for more detail.)
Table 1 summarizes the evaluation results of Llama 3.1-8B-Instruct and Llama 3.2-3B-Instruct on the English subset of LongBench with 2K, 4K, 6K, and 8K cache budgets using various eviction policies with block prompt processing enabled with $B=128$. As shown in Table 1, KeyDiff outperforms other eviction strategies across most tasks, even demonstrating better performance with smaller cache budgets. KeyDiff shows significant a improvement on the PassageRetrieval-en (PR-en) dataset, which tests whether long-term dependencies within a long prompt can be correctly recognized [4], while achieving near full-context model performance even with the smallest budget. Adding a sliding window to KeyDiff improves coding task performance (Table 13). We observed similar trends in the full LongBench task suite as shown in Table 10 and in the additional results in Appendix F.
KeyDiff exhibits similar or better performance compared to competing methods. Notably, the attention-based methods (e.g., H2O, TOVA, and SnapKV) show significant performance improvements over the $B=128$ case. This result supports our hypothesis: an eviction scheme robust to changes in the scope of comparison among tokens is essential in memory constrained environments where token-wise attention weight can’t be fully materialized.
Additional Results
We present the full evaluation results in Table 10 and more complete comparisons on LongBench in Appendix F, such as: standard prompt processing with a single large block (i.e. $B=\infty$) in Table 9; eviction method performance with Qwen 2.5-3B/7B-Instruct in Table 11; performance behavior with block sizes $B=[64,256]$ in Table 12; and performance on KeyDiff combined with a sliding window as described in Section 3.2. We also compare against the $L_{2}$-norm minimizing eviction method of [9] in Table 10.
4.3 Math-500 Reasoning Benchmark
Reasoning is an important long-context task for LLMs. Unlike other long-context use cases, reasoning typically involves a relatively short prompt followed by a long generation, which presents unique challenges for token eviction methods. To evaluate the effectiveness of token eviction methods, we apply KeyDiff and SnapKV to the DeepSeek-R1-Distill-Qwen-7B and Llama-8B distilled models [12], and assess their performance on the Math-500 reasoning benchmark [13]. Surprisingly, we found that Llama equipped with KeyDiff and a moderate KV cache budget performs comparably to, or slightly better than, the eviction-free baseline, while also outperforming SnapKV. We kindly refer the reader to Appendix E for additional details on the reasoning task evaluation.
@@ 4.4 Ablation Study
We evaluate the design choices of KeyDiff, including the similarity metrics and the choice of the anchor vector, and validate the efficacy of KeyDiff. We provide a full description of the test setup in Appendix G and summarize the findings here:
  • KeyDiff anchor choice does not greatly impact benchmark accuracies (See Table 14).
  • KeyDiff using cosine similarity as the distance metric outperforms other metrics (See Table 15)
4.5 Latency and Complexity
Additionally, in order to demonstrate that KeyDiff decreases end-to-end inference latency, we measured time to first token for the Llama 3.2 3B instruct model using different block prompt processing sizes and cache strategies. These results are visualized in Figures 17 and 7. Since KeyDiff does not require attention weight materialization, FlashAttention [8] can be used, resulting in up to 30% lower latency than TOVA and SnapKV. We compare the complexity of KeyDiff with competitors in Appendix B and perform a complete FLOP count in Section B.1.
Figure 7: Time-to-first-token (TTFT) for Llama 3.2-3B using Flash Attention with different eviction strategies with block prompt processing sizes $64$, $128$, and $256$.
@@ 5 Related Work
LLMs often exhibit high attention sparsity, where a small subset of keys receives a significant proportion of attention scores. This characteristic allows sparse approximation techniques to reduce the computational cost of attention. Similar to PagedAttention [18], Tang et al. [30] estimates the importance of a page (a contiguous set of keys) to a given query, whereas Rehg [26] further refined the budgets in a per-head manner. On the contrary, sample-based methods [44, 27] attempt to approximate token importance by inspecting the attention scores from the last few queries or certain query channel dimensions. Despite their effectiveness in reducing computational costs, these methods do not address the memory overhead of the KV cache, which typically retains all KVs.
@@
Different approaches to compress the KV cache include architecture modification such as GQA [2], which shares a KV cache across a small number of heads. Other techniques to compress the KV cache include quantization such as in [15, 23, 42] in which the authors use various techniques to take advantage of existing patterns to efficiently quantize and compress the KV cache. More related to our work [40] uses a scoring mechanism to determine the precision of the quantization for different tokens.
@@ 6 Conclusion
Inspired by our observation that distinctive keys tend to have high attention scores, we propose KeyDiff, a training-free KV cache eviction method based on key similarity that enables large language models to operate in memory and compute constrained environments. We justify KeyDiff by showing that it minimizes the pairwise cosine similarity among keys in the KV cache, maximizing the aforementioned diversity. KeyDiff significantly outperforms state-of-the-art KV cache eviction methods under similar memory constraints, with only a 1.5% and 0.04% accuracy drop from the non-evicting baseline while achieving 33% and 23% KV cache memory reduction on LongBench. Similar to other token eviction methods, KeyDiff is primarily designed and evaluated for the GQA attention mechanism used in models such as Llama and Qwen. In future work, we plan to extend KeyDiff for seamless integration with other attention variants, such as Multi-Head Latent Attention [12].
@@ A.1 Attention-based eviction methods
  • TOVA [24]: $\phi^{\mathrm{TOVA}}(A)=A_{-1,:}$,
  • H2O [43]: $\phi^{\mathrm{H2O}}(A)=A_{\mathrm{prev}}+A^{\top}1$,
  • SnapKV [35]: $\phi^{\mathrm{SnapKV}}(A)=(A^{\top}1)\ast K$, where $K$ is a vector of $\frac{1}{k}$ and $k$ is the kernel size of average smoothing.
Appendix B Runtime and Memory Complexity
@@
B.1 FLOP count of KeyDiff
The bulk of the computation in KeyDiff (neglecting the ‘topk‘ operator) is the following two expressions:
@@ B.1 FLOP count of KeyDiff
To compute the cosine similarity score, we have from above that $\mu(\hat{K})\cdot k_{i}$ requires $nd$ multiplications and $n(d-1)$ additions. Also from above, we have that computing $\|\mu(\hat{K})\|$ requires $d$ multiplications, $d-1$ additions and one square root. We reuse the computation of $\|k_{i}\|$ from the previous step and compute $\|k_{i}\|\|\mu(\hat{K})\|$ in $n$ multiplications and $\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)$ with more $n$ additions (assuming boolean comparison equals addition in cost). We can then divide through to compute $\frac{\mu(\hat{K})\cdot k_{i}}{\max(\|k_{i}\|\|\mu(\hat{K})\|,\,\epsilon)}$ with $n$ divisions. Therefore, computing the cosine similarity between the anchor and each key requires $nd+d+n$ multiplications, $nd+d-1$ additions, $n+1$ divisions and one square root.
@@
If, based on x86 instruction tables, we declare additions and square roots cost one FLOP (i.e. can be computed in one cycle), multiplications cost three FLOPs, division is roughly 47 FLOPs, we arrive at a final FLOP count of:
$943(3nd+d+n)+(3nd-n-1)+47*(2n+2)+(n+1)=(12d+97)n+3d+$ 94 . (A.1)
@@
Appendix C KeyDiff: A Theoretical Perspective
C.1 Additional PCA Visualizations
@@
(a) Layer 0 Head 1
(b) Layer 10 Head 1
(c) Layer 20 Head 1
(d) Layer 25 Head 1
Figure 11: PCA plots of Query and Keys from Llama-3.2-3B-Instruct
C.2 Derivation of KeyDiff from an Optimization Perspective
[…]
This is a combinatorial optimization problem, which is difficult to solve efficiently, particularly during inference. However, we can relax Section C.2 to produce a more tractable approximation to the original problem. First, we rewrite Section C.2 by normalizing keys $k_{i}$ such that $\hat{k}_{i}=\frac{k_{i}}{\|k_{i}\|}$ resulting in:
[…]
Key Cache Diversity and KeyDiff
To empirically verify KeyDiff’s ability to retain diverse keys, we apply PCA to the keys computed in an attention block of Llama 3.2-3B-Instruct after evaluating a long context prompt. We visualize the distribution of retained and evicted keys by applying sink attention [35], TOVA [24], and KeyDiff as eviction policies in Figure 11. Visual observation reveals the tokens retained by sink attention and TOVA tend to tightly cluster together while KeyDiff’s retained tokens are more evenly distributed. This observation generalizes across heads and layers, as shown in Section C.1.
We also visualize the log determinant of the Gram matrix of the key cache $\log(\det(KK^{T}))$ generated using different eviction policies in Figure 12. This quantity corresponds to the volume of space spanned by the keys in $\mathcal{C}$. The distribution of volumes for KeyDiff attains higher values, indicating that the retained keys are more distinctive than TOVA and sink, corroborating the results of Figure 11. Details of how these plots were generated are discussed in Section F.1.
C.3 Attention Sinks and Approximate Collinearity
(a) Keys
(b) Queries
Figure 13: Cosine similarity between keys and their mean (left) and queries and their mean (right) across heads and layers
Figure 14: Cosine similarity of mean key and mean query for each head and layer.
To better understand why there exists a negative correlation between cosine similarity between keys and attention scores, we look to recent research that seeks to the importance of attention sinks in decoder-based LLMs. The authors in [5] show that attention sinks emerge from training decoder-based LLMs since they can denoise the model and prevent rank collapse by limiting over mixing in attention heads. Moreover, attention patterns in decoder based models demonstrate that most attention logits are quite small (and almost always negative) for most keys and queries. This allows the attention sink to have high attention activation, preventing over mixing in addition to allowing the heads to specialize and selectively identify important tokens.
At the same time, the results in [11] suggest that hidden states, keys and queries are all approximately collinear in the sense that $\textrm{CosSim}(x_{i},x_{j})\gg 0$. In geometric terms, this means that most key tokens and query tokens lie in the same direction in Euclidean space. Our own results, seen in Figure 13 demonstrate that this is the case for both keys and queries. Our experiments show that most keys and queries lie within a small angular distance from the mean key and query. More than this, we see that across all heads, the mean key and mean query have negative cosine similarity Figure 14. Moreover, as seen in Figure 15, we find that the norms of keys and queries are tightly clustered around a relatively fixed value. This means that variations in the norm of key and query tokens have less impact on the magnitude of attention scores than their direction. These three observations indicate that most keys and queries combine to create uniformly small attention logits, and that larger attention weights are constructed by projecting keys closer to the direction of the mean query. This appears to be the fundamental mechanism through which over mixing is prevented: if most attention activations are very small, each head can increase the activations of a small number of keys across most queries selectively projecting them to be more aligned with the distribution of queries. This hypothesis is further supported by the fact that sink tokens themselves often have small norm, which results in an approximate no-op in the attention head as in [5], however in this case, the only way for a key corresponding to a sink token to have high attention scores is if it is as parallel as possible to the set of query tokens.
Figure 15: Distribution of $L^{2}$ norms for keys and queries across heads and layers.
Figure 16: Cosine similarity between highest KeyDiff scoring key token with mean query.
  • Most keys and queries have a similar L2 norm.
Theorem C.1.
Suppose that for a fixed query token $q$, there is a set of key tokens $\{k_{i}\}_{i=1}^{n}$ such that $||k_{i}||_{2}^{2}<M,\;\forall\;i$. Without loss of generality suppose $||q||=1$, the scaling parameter is $1$ and assume ${k^{*}}$ is a key not in $\{k_{i}\}_{i=1}^{n}$ with $||{k^{*}}||_{2}^{2}<M$ that has attention weight $w>0$:
Then
$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1\leq\textrm{CosSim}({k^{*}},q)$
Proof.
To show this we have that
$w\sum\limits_{i=1}^{n}\exp(k_{i}^{\top}q)$ $=(1-w)\exp({k^{*}}^{\top}q)$
Note that $-M\leq k_{i}^{\top}q\leq M$ and hence
$w\sum\limits_{i=1}^{n}\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
$wn\exp(-M)$ $\leq(1-w)\exp({k^{*}}^{\top}q)$
$\frac{wn\exp(-M)}{1-w}$ $\leq\exp({k^{*}}^{\top}q)$
$\ \log(wn)-M-\log(1-w)$ $\leq M\textrm{CosSim}({k^{*}},q)$
$\frac{\log(wn)-M-\log(1-w)}{M}$ $\leq\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{n}{n+1})-2M-\log(1-w)}{M}$ $\leq 2\textrm{CosSim}({k^{*}},q)$
$\frac{\log(\frac{n}{n+1})-\log(1-w)}{2M}-1$ $\leq\textrm{CosSim}({k^{*}},q)$
Taking the limit as $n\to\infty$ produces Lemma 3.1
In order to show this, we need the following auxiliary result.
Lemma C.2.
Proof.
$=\sum\limits_{i=1}^{n}\alpha_{i}x_{i}$
$=\sum\limits_{i=1}^{n}\langle\alpha_{i}x_{i},\alpha_{i}x_{i}\rangle$
$=\sum\limits_{i=1}^{n}\alpha_{i}^{2}$
proving the result. ∎
Theorem C.3.
Proof.
Consider the cosine similarity of $\bar{k}$ and ${k^{*}}$:
$\textrm{CosSim}(\bar{k},{k^{*}})$ $=\frac{{k^{*}}^{\top}\bar{k}}{||{k^{*}}||||\bar{k}||}$
$=\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}\alpha_{i}\beta_{i}$
$\leq\alpha_{q}\beta_{q}+\sum\limits_{i=1}^{n-1}|\alpha_{i}||\beta_{i}|$
Applying Young’s inequality we obtain
$=\alpha_{q}\beta_{q}+\frac{1}{2}(1-\alpha_{q}^{2})+\frac{1}{2}(1-\beta_{q}^{2})$
$=1+\alpha_{q}\beta_{q}-\frac{1}{2}\alpha_{q}^{2}-\frac{1}{2}\beta_{q}^{2}$
Appendix D TTFT Analysis
We have measured end-to-end inference latency (measured as time to first token (TTFT)) for Llama 3.2-3B via the standard Huggingface API when the model is using eager attention and FlashAttention as in Figure 17 and Figure 7. Tests are performed on NVIDIA A100 80GB GPUs. We test various block sizes and KV cache budgets. KeyDiff outperforms TOVA and SnapKV with FlashAttention as well as with eager attention for large cache budgets. We can see inference latency with KeyDiff is independent of block size when using FlashAttention because of KeyDiff’s linear complexity and its lack of required materialized attention weights.
Figure 17: TTFT for eager attention with different cache eviction strategies using block size $64$, $128$, $256$ for block prompt processing.
Appendix E Math-500 reasoning benchmark
In order to measure the effectiveness of different caching methods on reasoning tasks, we tested several different model using various caching algorithms on the Math 500 reasoning benchmark. Specifically, we test KeyDiff with a sliding window whose window size is $20\%$ of the KV cache budget, and SnapKV on the DeepSeek-R1-Distill-Llama-8B and DeepSeek-R1-Distill-Qwen-7B models. We randomly sample 5 responses with $\textrm{TopP}=0.95$, $\textrm{Temperature}=0.95$ with the 4096, 6144, and 32,786 max generation lengths. The reported scores are the average of accuracies over the random samples.
Table 3 summarizes the Math-500 evaluation results for Llama-8B. As shown in the table, token eviction methods generally perform well even with KV cache budgets that are strictly smaller than the maximum generation length. Surprisingly, we found that KeyDiff slightly outperforms baseline methods in certain configurations (e.g., KeyDiff with a 2K budget for 4K generation length, and KeyDiff with a 4K budget for 8K generation length).
To further analyze the effectiveness of token eviction, we measure accuracy in cases where the context length of the baseline method (i.e., prompt length + generation length) exceeds the available KV cache budget. As shown in Table 4, for samples where eviction is actively triggered, KeyDiff continues to outperform the token eviction baseline (SnapKV), and often achieves accuracy close to or slightly better than the non-evicting baseline.
We also conducted a similar evaluation on DeepSeek-R1-Distill-Qwen-7B and observed a slight performance degradation for token eviction methods compared to full KV cache baselines (See Tables 5 and 6.) However, KeyDiff still demonstrates comparable performance to SnapKV overall. This discrepancy may stem from architectural differences that Llama uses a lower GQA [2] ratio than Qwen, which results in more information compression in the KV cache. We hypothesize that models with more compression like Qwen are more sensitive to eviction since each evicted token contains more information in Qwen than Llama by design.
Table 3: Math 500 results on DeepSeek-R1-Llama-8B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
Method Max Gen. Length Budget Flex Match Exact Match Avg. Gen. Length
Full 4K N/A 0.711 0.537 2769
KeyDiff 4K 1024 0.695 0.531 2753
2048 0.720 0.546 2740
SnapKV 4K 1024 0.689 0.529 2759
2048 0.714 0.544 2757
Full 8K N/A 0.840 0.628 3812
KeyDiff 8K 2048 0.819 0.617 3888
4096 0.844 0.634 3805
SnapKV 8K 2048 0.805 0.610 3898
4096 0.828 0.627 3898
Full 32K N/A 0.898 0.668 6869
KeyDiff 32K 2048 0.883 0.662 7678
4096 0.894 0.668 7312
8192 0.894 0.667 7096
SnapKV 32K 2048 0.849 0.641 7509
4096 0.884 0.661 7218
8192 0.893 0.665 7005
Table 4: Math 500 results on DeepSeek-R1-Llama-8B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
Max Gen. Length $=$ 4K
Num Tokens $>$ 1K (497/500 samples)
Budget Flex Exact
Full N/A 0.709 0.534
KeyDiff 1024 0.693 0.528
SnapKV 1024 0.687 0.526
Num Tokens $>$ 2K (240/500 samples)
Full N/A 0.604 0.449
KeyDiff 2048 0.618 0.463
SnapKV 2048 0.610 0.458
Max Gen. Length $=$ 8K
Num Tokens $>$ 2K (353/500 samples)
Budget Flex Exact
Full N/A 0.783 0.584
KeyDiff 2048 0.756 0.570
SnapKV 2048 0.736 0.560
Num Tokens $>$ 4K (195/500 samples)
Full N/A 0.650 0.455
KeyDiff 4096 0.662 0.470
SnapKV 4096 0.637 0.453
Max Gen. Length $=$ 32K
Num Tokens $>$ 2K (353/500 samples)
Budget Flex Exact
Full N/A 0.783 0.584
KeyDiff 2048 0.756 0.570
SnapKV 2048 0.736 0.560
Num Tokens $>$ 4K (195/500 samples)
Full N/A 0.650 0.455
KeyDiff 4096 0.662 0.470
SnapKV 4096 0.637 0.453
Num Tokens $>$ 8K (162/500 samples)
Full N/A 0.793 0.550
KeyDiff 8192 0.786 0.545
SnapKV 8192 0.782 0.541
Table 5: Math 500 results on DeepSeek-R1-Qwen-7B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
Method Max Gen. Length Budget Flex Match Exact Match Avg. Gen. Length
Full 4K N/A 0.764 0.579 2630
KeyDiff 4K 1024 0.666 0.512 2692
2048 0.749 0.570 2629
SnapKV 4K 1024 0.692 0.533 2655
2048 0.749 0.566 2637
Full 8K N/A 0.877 0.658 3287
KeyDiff 8K 2048 0.811 0.613 3348
4096 0.867 0.647 3208
SnapKV 8K 2048 0.812 0.612 3328
4096 0.868 0.647 3214
Full 32K N/A 0.923 0.682 4051
KeyDiff 32K 2048 0.811 0.613 4322
4096 0.897 0.647 3800
8192 0.891 0.663 3741
SnapKV 32K 2048 0.812 0.612 4279
4096 0.868 0.647 3828
8192 0.891 0.662 3808
Table 6: Math 500 results on DeepSeek-R1-Qwen-7B distilled model (Higher is better). We highlight the methods showing the best performance within a given budget with boldface.
Max Gen. Length $=$ 4K
Num Tokens $>$ 1K (497/500 samples)
Budget Flex Exact
Full N/A 0.762 0.578
KeyDiff 1024 0.664 0.511
SnapKV 1024 0.690 0.532
Num Tokens $>$ 2K (328/500 samples)
Full N/A 0.650 0.478
KeyDiff 2048 0.628 0.464
SnapKV 2048 0.629 0.459
Max Gen. Length $=$ 8K
Num Tokens $>$ 2K (317/500 samples)
Budget Flex Exact
Full N/A 0.816 0.603
KeyDiff 2048 0.713 0.533
SnapKV 2048 0.718 0.533
Num Tokens $>$ 4K (179/500 samples)
Full N/A 0.626 0.441
KeyDiff 4096 0.611 0.419
SnapKV 4096 0.615 0.422
Max Gen. Length $=$ 32K
Num Tokens $>$ 2K (360/500 samples)
Budget Flex Exact
Full N/A 0.889 0.642
KeyDiff 2048 0.713 0.533
SnapKV 2048 0.718 0.533
Num Tokens $>$ 4K (141/500 samples)
Full N/A 0.787 0.524
KeyDiff 4096 0.611 0.419
SnapKV 4096 0.615 0.422
Num Tokens $>$ 8K (53/500 samples)
Full N/A 0.550 0.275
KeyDiff 8192 0.392 0.222
SnapKV 8192 0.381 0.215
Appendix F Additional discussion for LongBench
F.1 Empirical Motivation for KeyDiff Setup